calculate the number of photons emitted in 10 hr by a 60w sodium lamp of wavelength 5893 A

Grade:8

1 Answers

Arun
5 years ago
Given , Power = 60W
time = 10 hours = 10 × 3600 sec = 3.6 × 10⁴ sec 
We know, Energy = Power × time 
so, Energy = 60W × 3.6 × 10⁴ sec = 21.6 × 10⁵J 
Energy = 2.16 × 10⁶J 
Now, A/C planks theory, 
Energy of one photon = hc/λ , 
Here h is plank's constant and λ is wavelength of light emitted by sodium lamp.

it is given λ = 5893 A° = 5.893 × 10⁻⁷ m 
Now, Energy of one photon = 6.63 × 10⁻³⁴ × 3 × 10⁸/5.893 × 10⁻⁷ 
= 19.89/5.893 × 10⁻¹⁹ J 
Now, number of photons = energy /energy of one photon 
= 2.16 × 10⁶/{19.89/5.983 × 10⁻¹⁹} 
= {2.16 × 5.893/19.89 }× 10²⁵ 
= 0.64 × 10²⁵
number of photons = 6.4 × 10²⁴